OpenStudy (song_of_the_sole):

to the nearest millimeter, a cell phone is 76mm long and 30mm wide. what is the ratio of the width to the length?

OpenStudy (anonymous):

30 to 76 put it in proper terms

OpenStudy (song_of_the_sole):

how would i do that?

OpenStudy (anonymous):

30 is your width as stated. 76 is you length

OpenStudy (anonymous):

let me show you

OpenStudy (anonymous):

30:76

OpenStudy (mathstudent55):

A ratio is a comparison of two numbers by division. Since you are asked for the ratio of the width to the length, that means you divide the width by the length.

OpenStudy (anonymous):

30 mm in width to 76 mm in length

OpenStudy (anonymous):

Hope you understand.

OpenStudy (song_of_the_sole):

ait so do i just put 30:76?

OpenStudy (anonymous):

Oh, I use to be like you, thinking it was an equation

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

I mean, explain if it's not multiple choice

OpenStudy (song_of_the_sole):

but it says its wrong

OpenStudy (song_of_the_sole):

how would i put 30:76 into the ratios as a simplified fraction?

OpenStudy (mathstudent55):

A ratio can be expressed in several different ways. The ratio of x to y can be written as: a:y a/y a to y If the two numbers are divisible by a number, you can do the division. It's the same as writing a fraction in lowest terms.

OpenStudy (anonymous):

You didn't specify

OpenStudy (song_of_the_sole):

what do you mean?

OpenStudy (mathstudent55):

Since both 30 and 76 are even, what number are they both divisile by?

OpenStudy (song_of_the_sole):

2

OpenStudy (mathstudent55):

Good. Start by dividing both by 2.

OpenStudy (song_of_the_sole):

30/2=15

OpenStudy (mathstudent55):

good

OpenStudy (song_of_the_sole):

76/2=38

OpenStudy (mathstudent55):

also good Now you know you can write the ratio as 15:38

OpenStudy (mathstudent55):

Can you reduce it more?

OpenStudy (song_of_the_sole):

no because 15 is odd and 38 is even

OpenStudy (song_of_the_sole):

right?

OpenStudy (mathstudent55):

You are correct, but that is not the reason. For example, 18 and 21 are both divisible by 3 even though one is even and one is odd.

OpenStudy (mathstudent55):

What are the prime factors of 15? 15 = 3 * 5 What are the prime factors of 38? 38 = 2 * 19

OpenStudy (song_of_the_sole):

so is15:38 the ratio as a simplified fraction?

OpenStudy (mathstudent55):

The prime factors of 15 and 38 are all different numbers. There is no prime factor in common, so 15/38 is completely reduced.

OpenStudy (mathstudent55):

Yes. 15:38 or 15/38 or "15 to 38" are all the simplified ratio you need.

OpenStudy (song_of_the_sole):

oh okay thank you :) @mathstudent55